OpenStudy (anonymous):

check my answer ? x-2/x+4=x-6/x-4 I got x=8 as a result

OpenStudy (danjs):

is it right to assume each one of those is a quantity? (x-2) / (x+4) = (x-6)/(x-4)

OpenStudy (anonymous):

yes, sorry.

OpenStudy (danjs):

plug in 8 for x, is it true

OpenStudy (danjs):

6 / 12 = 2 / 4 1/2 = 1/2 right

OpenStudy (anonymous):

so im right?

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

yea, x=8 makes that equation true

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

The answer is X=-1/2